# MOOSE-Star: Unlocking Tractable Training for Scientific Discovery by Breaking the Complexity Barrier

### 04 Mar 2026 v1

@@ 1 Introduction
To support this data-hungry training paradigm, we developed a comprehensive pipeline that processes scientific literature across domains such as biology, chemistry, and cognitive science. In total, we extracted research backgrounds, ground-truth hypotheses, and inspirations—each explicitly linked to a historical citation—from 108,717 papers, a massive computational effort consuming approximately 38,400 A800 GPU hours. We release this processed dataset as the TOMATO-Star dataset.
@@
Our contributions can be summarized as below:
@@ 2 Related Work
While a significant body of research focuses on novel inference algorithms for LLM-based discovery (Romera-Paredes et al., 2024; Novikov et al., 2025; Shojaee et al., 2025; Lu et al., 2024), few studies investigate methods for training LLMs for discovery. Most existing training efforts focus on constructing external feedback mechanisms for generated hypotheses rather than modeling the core conditional probability $P(h|b)$ itself. For instance, Weng et al. (2024); Li et al. (2024) leverage peer review data to train critique models, which serve as reward signals for the discovery agent. Similarly, in the domain of equation discovery, Behzadifar et al. (2025) derive rewards based on the goodness of fit between the proposed hypothesis (equation) and observed data. Goel et al. (2025) extract evaluation criteria from scientific literature to construct rubric-based LLM self-evaluators, while Pu et al. (2025) train models to retrieve ancillary high-level guidance from a constrained set to assist discovery. O’Neill et al. (2025) attempt to model $P(h|b)$ directly; however, obtaining the reasoning trace towards $h$ from only $b$ via distillation is prohibitive (as discussed in § 6), and generating simulated reasoning traces directly from $b$ and the ground truth $h$ as input is very difficult to work (Wang et al., 2025).
This work focuses exclusively on training LLMs for scientific discovery, distinct from related fields such as general scientific reasoning (Bai et al., 2025) or modeling of specific science modality (Jumper et al., 2021).
@@ 3 Preliminary: The Decomposition Theory
This compositional view is well-supported by cognitive science literature (Koestler, 1964) and has proven empirically effective across scientific disciplines (Liu et al., 2025). For instance, the discovery of Backpropagation can be viewed as composing the background of “multilayer logistic regression” with the inspiration of the “chain rule” from calculus; similarly, a new synthesis method in chemistry might be derived by applying a specific “catalyst” (inspiration) to a known reaction mechanism.
@@ 4 Dataset Construction
Data Collection and Split. We collected 108,717 open-access papers from the NCBI database (Schoch et al., 2020), spanning biology, chemistry, and cognitive science. The corpus covers the period from January 2020 to October 2025. To rigorously evaluate generalization, we enforce a strict temporal split: papers published from January 2020 to September 2025 constitute the training set, while papers from October 2025 serve as the held-out test set. We also perform explicit checking to make sure the two sets do not have overlap with each other. October 2025 is latest enough, so that the evaluation results should not be contaminated.
@@ 5.2 Method $\mathrm{I}$: Decomposed Sequential Training ($O(N)$)
For the IR task, the model is trained to identify the target inspiration via a generative selection process. The input consists of the query context and a candidate pool of 15 papers (1 positive and 14 negatives), each represented by its title and abstract. The negative set comprises easy negatives (randomly sampled papers published no later than the source paper) and hard negatives drawn from three specific categories: papers with keyword overlap with the source background, papers with keyword overlap with the ground-truth inspiration, and papers semantically similar to the ground-truth inspiration. We acknowledge that these negatives might occasionally indicate valid alternative inspirations. However, the ground-truth citation represents the author’s explicit acknowledgment of causal influence, providing the most reliable supervision signal for prioritizing optimal inspirations over merely related ones. Furthermore, this single-selection objective minimizes the penalty for potentially valid unlabelled candidates compared to independently labeling each candidate.
For the HC task, the objective is to generate the incremental hypothesis update, $\Delta h_{j}$. The model is conditioned on the research background $b$, the retrieved ground-truth inspiration $i_{j}$, and $h_{j-1}$, represented by the concatenation of $\{\Delta h_{0},\ldots,\Delta h_{j-1}\}$ (if $j\geq 1$). To enable effective rejection sampling, we developed a rubric-based evaluator to compare the quality of sampled hypotheses against the ground-truth $\Delta h_{j}$, ensuring that only high-quality samples are retained for training. The evaluation rubric is provided in § C.
@@ 5.3 Method $\mathrm{II}$: Bounded Composition
We achieve this by relaxing the retrieval objective. In § 5.2, the formulation assumes a strict “Exact Match” constraint: the IR module must identify the unique ground-truth inspiration $i^{*}$ (11 out of $N$), and the HC module relies solely on this exact input. We propose Bounded Composition, which introduces a Semantic Tolerance Space of size $M$. Instead of requiring $i^{*}$, we posit that the HC module can be trained to function robustly within a semantic neighborhood $\mathcal{I}_{i^{*}}$ centered on $i^{*}$ containing $M$ candidates (where $i^{*}\in\mathcal{I}_{i^{*}}\subset\mathcal{I}$). If the retriever provides any inspiration $i\in\mathcal{I}_{i^{*}}$, the HC module is expected to reason through this proxy $i$ to recover the underlying $i^{*}$ and correctly compose the hypothesis.
This relaxation alters the complexity landscape:
@@
To operationalize Bounded Composition, we first curate a pool of semantic proxies. For each ground-truth inspiration $i^{*}$ in the dataset, we query Semantic Scholar to retrieve up to 50 semantically similar papers. We embed these candidates using SPECTER2 (Singh et al., 2022) and compute their cosine similarity to $i^{*}$. Based on these scores, we stratify valid candidates into three difficulty tiers: Easy ($[0.94,0.97)$), Medium ($[0.92,0.94)$), and Hard ($[0.90,0.92)$). From each available tier, we select the single candidate with the highest similarity score to serve as the tier representative. We then employ the teacher model to generate hypotheses conditioned on these proxy inspirations rather than the exact ground truth. The generated outputs are rigorously evaluated against the quality control rubric described in § C. Crucially, for instances where proxies from multiple tiers yield passing hypotheses, we prioritize the sample from the most difficult tier (i.e., lowest similarity) for the final training set. This strategy explicitly maximizes the learned semantic tolerance space, forcing the model to reason robustly even with less precise retrieval contexts.
@@ 5.4 Method $\mathrm{III}$: Hierarchical Search ($O(\log N)$)
Despite the reduction to $O(N/M)$, a linear scan over the global knowledge base $\mathcal{I}$ remains inefficient. To enable logarithmic-time access (at least in the best-case scenario), we propose replacing the flat scan with a Hierarchical Best-First Search. This design draws inspiration from MOOSE-Chem2 (Yang et al., 2025a), which formulates fine-grained hypothesis discovery as a hierarchical search over varying levels of detail abstraction. Here, we adapt this concept by constructing the hierarchy directly over the inspiration space $\mathcal{I}$. By organizing $\mathcal{I}$ into a semantic search tree, we can rapidly navigate to the most promising subspaces, pruning irrelevant branches early in the process.
@@
During inference, we traverse the tree using a Best-First Search (BFS) strategy. We maintain a priority queue of candidate nodes, ranked by a path confidence score. In each iteration, we pop the highest-scoring node for expansion. The IR model evaluates the sub-branches of this node, generating a probability distribution over its children (derived from the output token log-probabilities). Crucially, these local probabilities are aggregated with the parent node’s stored cumulative probability to compute updated scores for each child, which are then pushed back into the priority queue. To ensure fair comparison between nodes at different depths—avoiding the bias against deeper paths inherent in raw cumulative products—we define the ranking score as the geometric mean of the probabilities along the path from the root:
@@ 6.1 Method $\mathrm{I}$: Decomposed Sequential Training
R1-Distilled-Qwen-7B 4.34 1.96 1.40 0.97 231.02
MS-HC-7B 5.08 2.21 1.58 1.29 204.12
   w/ $1\times$ bounded 5.16 2.24 1.60 1.31 203.84
   w/ $2\times$ bounded 5.15 2.25 1.58 1.32 205.17
Inspiration Retrieval (IR).
We fine-tuned R1-Distilled-Qwen-7B on a dataset of 152k retrieval examples, resulting in MS-IR-7B (where MS stands for MOOSE-Star). As shown in Table 1, this specialized training yields a substantial performance gain, boosting accuracy from 28.42% to 54.37%.
Hypothesis Composition (HC).
We fine-tuned R1-Distilled-Qwen-7B on 97k samples generated via teacher-based rejection sampling, resulting in MS-HC-7B (MS stands for MOOSE-Star). Performance is evaluated using the M3 score (detailed in § C), an LLM-based rubric summing scores for Motivation, Mechanism, and Methodology. Each dimension ranges from [0, 4] (representing 0%–100% information recall), yielding a total score of 12.
@@
R1-Distilled-Qwen-7B 2.58 1.43 0.62 0.54 2.16 1.24 0.46 0.45 1.82 1.13 0.33 0.36
MS-HC-7B 2.82 1.53 0.62 0.67 2.37 1.36 0.47 0.55 2.20 1.30 0.40 0.49
   w/ $1\times$ bounded 2.96 1.55 0.70 0.71 2.52 1.41 0.53 0.58 2.19 1.25 0.44 0.49
   w/ $2\times$ bounded 3.03 1.62 0.69 0.72 2.49 1.37 0.53 0.59 2.25 1.29 0.44 0.52
6.2 Method $\mathrm{II}$: Bounded Composition
Table 3 evaluates performance under noisy conditions, using proxy inspirations stratified by semantic similarity to $i^{*}$ (Easy, Medium, Hard). The results demonstrate that incorporating bounded training data significantly boosts M3 scores across all tiers compared to the base model.
6.3 Method $\mathrm{III}$: Hierarchical Search
@@ 6.4 Method $\mathrm{IV}$: Motivation Planning
7 Scaling Analysis
7.1 Scaling Brute-force Sampling $P(h|b)$
Table 5: Comparison of sampling pass rates between Brute-force (End-to-End) and Hypothesis Composition (Decomposed). While Brute-force performance collapses to 0.00% as complexity ($k$) increases, the decomposed HC module maintains a robust 47.33% pass rate, unlocking the data engine for RFT.
@@ 7 Scaling Analysis
HC $P(\Delta h|i,b)$ 1 (per step) 195,301 92,427 47.33%
Direct distillation from the end-to-end distribution $P(h|b)$ is intractable due to the extreme scarcity of valid reasoning traces. Table 5 illustrates this bottleneck: finding a qualified hypothesis is rare even for simple cases ($k=1$), and becomes statistically impossible for multi-inspiration trajectories ($k\geq 2$) due to combinatorial explosion.
In contrast, the decomposed HC step demonstrates a high success rate (47.33%). This performance gain is driven by two structural advantages. Primarily, by isolating a single inspiration step, the HC task circumvents the combinatorial complexity of the full problem: it is only required to solve for one incremental innovation $\Delta h_{j}$ rather than the joint distribution of all $k$ innovations. Secondarily, the decomposition outsources the search burden to the IR module; thus, conditioning on the provided inspiration $i^{*}$ explicitly guides the generation. Together, these factors empirically validate the complexity analysis in § 5.
@@
To understand why this performance gap determines training feasibility, we must examine the criteria used for harvesting RFT training data (detailed in § C). Both strategies use the same rubric, evaluating a generated hypothesis component against the ground truth across Motivation, Mechanism, and Methodology. To serve as a positive sample for RFT, we impose a filtering threshold: the generation must achieve a Total Score $\geq 8$ (out of 12).
The structural difference lies in how this threshold applies to the two paradigms:
  • Brute-force: Since the baseline generates the hypothesis $h$ holistically, a valid sample for a paper with $k\geq 2$ must successfully meet the filtering threshold for every constituent innovation $\Delta h_{j}^{*}$ ($j=1\dots k$) simultaneously. This strict conjunctive constraint drives the success rate to near-zero, creating a “training deadlock”: the model cannot generate enough valid samples to initiate RFT, rendering the training process impossible.
  • HC: In contrast, the HC module is trained to synthesize a single increment $\Delta h_{j}$. Crucially, leveraging the decomposed data construction, we can condition the model on the ground-truth inspiration $i_{j}^{*}$. Under this setup, the model only needs to meet the threshold for the corresponding single $\Delta h_{j}^{*}$. Consequently, this efficient sampling capability breaks the training deadlock, making it feasible to train HC using RFT.

### 06 Apr 2026 v2

@@ 1 Introduction
To support this data-hungry training paradigm, we developed a comprehensive pipeline that processes scientific literature across domains such as biology, chemistry, medicine, medical imaging, psychology, and cognitive science. In total, we extracted research backgrounds, ground-truth hypotheses, and inspirations—each explicitly linked to a historical citation—from 108,717 papers, a massive computational effort consuming approximately 38,400 A800 GPU hours. We release this processed dataset as the TOMATO-Star dataset.
@@
Our contributions can be summarized as follows:
@@ 2 Related Work
While a significant body of research focuses on novel inference algorithms for LLM-based discovery (Romera-Paredes et al., 2024; Novikov et al., 2025; Shojaee et al., 2025; Lu et al., 2024), few studies investigate methods for training LLMs for discovery. Most existing training efforts focus on constructing external feedback mechanisms for generated hypotheses rather than modeling the core conditional probability $P(h|b)$ itself. For instance, Weng et al. (2024); Li et al. (2024) leverage peer review data to train critique models, which serve as reward signals for the discovery agent. Similarly, in the domain of equation discovery, Behzadifar et al. (2025) derive rewards based on the goodness of fit between the proposed hypothesis (equation) and observed data. Goel et al. (2025) extract evaluation criteria from scientific literature to construct rubric-based LLM self-evaluators, while Pu et al. (2025) train models to retrieve ancillary high-level guidance from a constrained set to assist discovery. O’Neill et al. (2025) attempt to model $P(h|b)$ directly; however, obtaining the reasoning trace towards $h$ from only $b$ via distillation is prohibitive (as discussed in § 7.1), and simulated reasoning traces generated directly from $b$ and the ground-truth $h$ are often difficult to make effective for training (Wang et al., 2025).
This work focuses exclusively on LLMs for scientific discovery, distinct from related fields such as general scientific reasoning (Bai et al., 2025) or modeling of specific science modality (Jumper et al., 2021).
@@ 3 Preliminary: The Decomposition Theory
This compositional view is well-supported by cognitive science literature (Koestler, 1964) and has proven empirically effective across scientific disciplines (Liu et al., 2025). For instance, the discovery of Backpropagation can be viewed as composing the background of “multilayer logistic regression” with the inspiration of the “chain rule” from calculus; similarly, a new synthesis method in chemistry might be derived by applying a specific “catalyst” (inspiration) to a known reaction mechanism (background).
@@ 4 Dataset Construction
Data Collection and Split. We collected 108,717 open-access papers from the NCBI database (Schoch et al., 2020), spanning biology, chemistry, medicine, medical imaging, psychology, and cognitive science. The corpus covers the period from January 2020 to October 2025. To rigorously evaluate generalization, we enforce a strict temporal split: papers published from January 2020 to September 2025 constitute the training set, while papers from October 2025 serve as the held-out test set. We also perform explicit checking to make sure the two sets do not have overlap with each other. October 2025 is latest enough, so that the evaluation results should not be contaminated.
@@ 5.2 Method $\mathrm{I}$: Decomposed Sequential Training ($O(N)$)
For the IR task, the model is trained to identify the target inspiration via a generative selection process. The input consists of the query context and a pool of 15 papers (1 positive and 14 negatives), each represented by its title and abstract. The negative set comprises easy negatives (randomly sampled papers published no later than the source paper) and hard negatives drawn from three categories: papers with keyword overlap with the source background, papers with keyword overlap with the ground-truth inspiration, and papers semantically similar to the ground-truth inspiration. We acknowledge that these negatives might occasionally indicate valid alternative inspirations. However, the ground-truth citation represents the author’s explicit acknowledgment of causal influence, providing the most reliable supervision for prioritizing optimal inspirations over merely related ones. Furthermore, this single-selection objective minimizes the penalty for potentially valid unlabelled candidates compared to independently labeling each candidate.
For the HC task, the objective is to generate the incremental hypothesis update, $\Delta h_{j}$. The model is conditioned on the research background $b$, the retrieved ground-truth inspiration $i_{j}$, and $h_{j-1}$, represented by the concatenation of $\{\Delta h_{0},\ldots,\Delta h_{j-1}\}$ (if $j\geq 1$). To enable effective rejection sampling, we developed a rubric-based evaluator to compare the quality of sampled hypotheses against the ground-truth $\Delta h_{j}^{*}$, ensuring that only high-quality samples are retained for training. The evaluation rubric is provided in § E.
@@ 5.3 Method $\mathrm{II}$: Bounded Composition
We achieve this by relaxing the retrieval objective. In § 5.2, the formulation assumes a strict “Exact Match” constraint: the IR module must identify the unique ground-truth inspiration $i^{*}$ (11 out of $N$), and the HC module relies solely on this exact input. We propose Bounded Composition, which introduces a Semantic Tolerance Space of size $M$. Instead of requiring $i^{*}$, we posit that the HC module can be trained to function robustly within a semantic neighborhood $\mathcal{I}_{i^{*}}$ centered on $i^{*}$ containing $M$ candidates (where $i^{*}\in\mathcal{I}_{i^{*}}\subset\mathcal{I}$). If the retriever provides any inspiration $i\in\mathcal{I}_{i^{*}}$, the HC module is expected to reason through this proxy $i$ to recover the underlying $i^{*}$ and correctly compose the hypothesis. This relaxation alters the complexity landscape:
@@
To operationalize Bounded Composition, we first curate a pool of semantic proxies. For each ground-truth inspiration $i^{*}$ in the dataset, we query Semantic Scholar to retrieve up to 50 semantically similar papers. We embed these candidates using SPECTER2 (Singh et al., 2022) and compute their cosine similarity to $i^{*}$. Based on these scores, we stratify valid candidates into three difficulty tiers: Easy ($[0.94,0.97)$), Medium ($[0.92,0.94)$), and Hard ($[0.90,0.92)$). From each available tier, we select the single candidate with the highest similarity score to serve as the tier representative. We then employ the teacher model to generate hypotheses conditioned on these proxy inspirations rather than the exact ground truth. The generated outputs are rigorously evaluated against the quality control rubric described in § E. Crucially, for instances where proxies from multiple tiers yield passing hypotheses, we prioritize the sample from the most difficult tier (i.e., lowest similarity) for the final training set. This strategy explicitly maximizes the learned semantic tolerance space, forcing the model to reason robustly even with less precise retrieval contexts.
@@ 5.4 Method $\mathrm{III}$: Hierarchical Search ($O(\log N)$)
Despite the reduction to $O(N/M)$, a linear scan over the global knowledge base $\mathcal{I}$ remains inefficient. To enable logarithmic-time search (at least in the best-case scenario), we replace the flat scan with a Hierarchical Best-First Search. This design is inspired by MOOSE-Chem2 (Yang et al., 2025a), which formally introduces the fine-grained hypothesis discovery task, formulates it as an optimization problem, and demonstrates that this optimization process can be made more tractable by organizing it hierarchically over multiple abstraction levels. While MOOSE-Chem2 develops this principle in the hypothesis space, here we adapt it to inspiration retrieval by constructing the hierarchy directly over the inspiration space $\mathcal{I}$.
@@
During inference, we traverse the tree using a Best-First Search strategy. We maintain a priority queue of candidate nodes, ranked by a path confidence score. In each iteration, we pop the highest-scoring node for expansion. The IR model evaluates the sub-branches of this node, generating a probability distribution over its children (derived from the output token log-probabilities). Crucially, these local probabilities are aggregated with the parent node’s stored cumulative probability to compute updated scores for each child, which are then pushed back into the priority queue. To ensure fair comparison between nodes at different depths—avoiding the bias against deeper paths inherent in raw cumulative products—we define the ranking score as the geometric mean of the probabilities along the path from the root:
@@ 6.1 Method $\mathrm{I}$: Decomposed Sequential Training
Frontier Models
GPT-4o-mini 41.99%
Claude-Sonnet-4.6 45.02%
DeepSeek-R1 45.11%
GPT-4o 48.37%
Gemini-3-Flash 51.44%
GPT-5.4 51.50%
Gemini-3-Pro 54.89%
Baselines
@@
Frontier Models
GPT-4o-mini 5.18 2.26 1.64 1.28 175.29
GPT-4o 5.99 2.43 1.95 1.60 264.93
DeepSeek-R1 6.42 2.54 2.10 1.77 264.46
Gemini-3-Flash 6.42 2.48 2.16 1.79 351.79
Gemini-3-Pro 6.70 2.59 2.26 1.85 300.86
Claude-Sonnet-4.6 7.28 2.78 2.41 2.08 510.26
GPT-5.4 7.82 2.99 2.61 2.22 504.60
Baselines
R1-Distilled-Qwen-7B 4.05 1.96 1.30 0.80 231.02
MS-HC-7B 4.68 2.13 1.46 1.09 204.12
   w/ $1\times$ bounded 4.74 2.16 1.48 1.10 203.84
   w/ $2\times$ bounded 4.73 2.15 1.48 1.09 205.17
MS-7B 5.02 2.22 1.59 1.20 208.98
@@
R1-Distilled-Qwen-7B 2.72 1.50 0.70 0.51 2.27 1.35 0.55 0.38 2.00 1.24 0.46 0.31
MS-HC-7B 3.10 1.63 0.81 0.66 2.59 1.46 0.61 0.53 2.39 1.38 0.54 0.48
   w/ $1\times$ bounded 3.09 1.61 0.81 0.67 2.72 1.48 0.66 0.57 2.42 1.39 0.55 0.47
   w/ $2\times$ bounded 3.18 1.67 0.83 0.68 2.74 1.51 0.66 0.57 2.56 1.45 0.58 0.53
MS-7B 3.37 1.72 0.90 0.75 2.86 1.55 0.69 0.62 2.78 1.52 0.66 0.60
Inspiration Retrieval (IR).
We fine-tuned R1-Distilled-Qwen-7B on a dataset of 152k retrieval examples, resulting in MS-IR-7B (where MS stands for MOOSE-Star). As shown in Table 1, this specialized training yields a substantial performance gain, boosting accuracy from 28.42% to 54.37%. We verify in § C that this improvement reflects genuine retrieval ability rather than exploitation of superficial patterns.
Hypothesis Composition (HC).
We fine-tuned R1-Distilled-Qwen-7B on 97k samples generated via teacher-based rejection sampling, resulting in MS-HC-7B (MS stands for MOOSE-Star). Performance is evaluated using the M3 score (detailed in § E), an LLM-based rubric summing scores for Motivation, Mechanism, and Methodology. Each dimension ranges from [0, 4] (representing 0%–100% information recall), yielding a total score of 12. We use GPT-4o (OpenAI, 2024) as the judge model.
Overall.
These results suggest an interesting asymmetry between the two subtasks. For inspiration retrieval, targeted post-training appears sufficient to elevate a 7B model to near-frontier performance. By contrast, although hypothesis composition improves substantially over the base model, it remains clearly below frontier LLMs. This pattern suggests that retrieval may rely more on activating and organizing latent scientific associations already present in the base model, whereas high-quality composition likely depends more strongly on the base model’s underlying scientific reasoning and generation capacity.
Motivated by the complementary nature of the two tasks, we jointly train on both IR and HC data (w/ $1\times$ bounded), yielding a unified model MS-7B. As shown in Tables 1 and 2, MS-7B retains IR accuracy (54.34% vs. 54.37%) while improving HC performance (5.02 vs. 4.68), suggesting that multi-task training provides a beneficial inductive bias for hypothesis composition without degrading retrieval.
@@
6.2 Method $\mathrm{II}$: Bounded Composition
Table 3 evaluates performance under noisy conditions, using proxy inspirations stratified by semantic similarity to $i^{*}$ (Easy, Medium, Hard). The results demonstrate that incorporating bounded training data significantly boosts M3 scores across all tiers compared to the base model. MS-7B further improves across all tiers, with the largest gain on Hard proxies (+0.22 total over w/ $2\times$ bounded), indicating that multi-task training strengthens robustness to retrieval noise. Similarly, GPT-4o is used as the judge model.
6.3 Method $\mathrm{III}$: Hierarchical Search
@@ 6.4 Method $\mathrm{IV}$: Motivation Planning
Table 5: Comparison of sampling pass rates between Brute-force (End-to-End) and Hypothesis Composition (Decomposed). Both generation and rubrics-based quality scoring are performed by R1-Distilled-Qwen-32B. While Brute-force collapses to 0.00% as complexity ($k$) increases, the decomposed HC module maintains a stable $\sim$47–52% per-step pass rate regardless of $k$, unlocking the data engine for RFT.
@@
[…]
Total 195,301 96,879 49.60%
7 Scaling Analysis
7.1 Scaling Brute-force Sampling $P(h|b)$
Direct distillation from the end-to-end distribution $P(h|b)$ is intractable due to the extreme scarcity of valid reasoning traces. Table 5 illustrates this bottleneck: using R1-Distilled-Qwen-32B as both the generator and the rubric-based scorer, finding a qualified hypothesis is rare even for simple cases ($k=1$), and becomes statistically impossible for multi-inspiration trajectories ($k\geq 2$) due to combinatorial explosion.
In contrast, the decomposed HC step demonstrates a high success rate (49.60%). This performance gain is driven by two structural advantages. Primarily, by isolating a single inspiration step, the HC task circumvents the combinatorial complexity of the full problem: it is only required to solve for one incremental innovation $\Delta h_{j}$ rather than the joint distribution of all $k$ innovations. Secondarily, the decomposition outsources the search burden to the IR module; thus, conditioning on the provided inspiration $i^{*}$ explicitly guides the generation. Together, these factors empirically validate the complexity analysis in § 5.
@@ 7 Scaling Analysis
Both strategies are scored by the same rubric (§ E), which evaluates Motivation, Mechanism, and Methodology against the ground truth. A generation must achieve a Total Score $\geq 8$ (out of 12) to qualify as a positive sample for RFT.
  • Brute-force: Since the baseline generates the hypothesis $h$ holistically, a valid sample for a paper with $k$ steps must meet the threshold for every constituent innovation $\Delta h_{j}^{*}$ ($j=1\dots k$) simultaneously. As $k$ increases, the likelihood of satisfying all $k$ criteria jointly vanishes, creating a “training deadlock” where the model cannot generate enough valid samples to initiate RFT.
  • HC: In contrast, the HC module synthesizes a single increment $\Delta h_{j}$ conditioned on the corresponding inspiration $i_{j}^{*}$. The reduced scope and provided guidance together yield a stable $\sim$50% pass rate, breaking the training deadlock and making RFT feasible.